# Enhancing ground-state interaction strength of neutral atoms via Floquet stroboscopic dynamics

### v1

@@ I Introduction
Ground-state neutral atoms exhibit long coherence times and excellent isolation from environmental noise, making them promising candidates for scalable quantum technologies [1, 2, 3, 4, 5, 6]. Meanwhile, coupling ground states to Rydberg levels offers a complementary advantage: strong, long-range interactions via van der Waals or dipole-dipole coupling, which enable fast entangling operations through the Rydberg blockade effect [7, 8, 9, 10, 11, 12]. The Rydberg blockade provides a powerful mechanism for engineering strong, controllable interactions and creating collective quantum states in which a single excitation is coherently shared throughout the atomic ensemble [13, 14, 15]. Such collective control has opened broad application prospects in quantum optics [16, 17, 18, 19, 20, 21, 22], quantum computing [23, 24, 25], quantum communication [26, 27, 28, 29, 30], and single-photon sources [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. With the development of experimental technology, the coherent manipulation of Rydberg atomic ensembles is driving a deeper understanding of collective quantum dynamics and enabling new opportunities for scalable and high-performance quantum technologies [42, 43, 44, 45, 46, 47, 48, 49, 50, 51].
@@
The finite lifetime and strong decoherence effects inherent to Rydberg states present significant challenges for high-fidelity quantum operations. This naturally raises the question: Can sufficiently strong and controllable interactions be engineered while atoms remain in their ground state? In this context, finding ways to achieve effective interactions without exciting high-energy states is especially important. Just as the Rydberg blockade allows only a single collective excitation in Rydberg states, strong interactions between ground states can also give rise to a ground-state blockade mechanism [52, 53, 54, 55, 56]. This single-excitation state is of great significance for photon storage [57, 58],single-photon sources [59, 60, 61, 62, 63], quantum state transfer [64, 65], and other quantum information technologies [66, 67]. Compared to the complexity and strong interactions inherent in Rydberg states, the single-excitation state of ground states provides a more precise and stable quantum platform, which makes it particularly promising for applications in quantum computing and quantum simulation [68, 69, 70, 71].
@@
In this work, we utilize Floquet stroboscopic dynamics to prepare a collective $W$ state within the ground-state manifold of a Rydberg atomic ensemble [72, 73, 74, 75, 76, 77, 78, 79, 80]. The system evolves from the collective ground state $|G\rangle=|g_{1}g_{2}...g_{\mathbb{N}}\rangle$ to the symmetric single-excitation state $|W_{\mathbb{N}}\rangle={1/\sqrt{\mathbb{N}}\textstyle\sum_{i=1}^{\mathbb{N}}}|g_{1}g_{2}...e_{i}...g_{\mathbb{N}}\rangle$, while the double excitations remain unpopulated at stroboscopic time points. This indirectly induces an effective interaction between ground-state neutral atoms, leading to a strong blockade effect that enables high-fidelity preparation of the $W$ state. Finally, we analyze the application of the scheme to single-photon generation. Our approach remains effective not only within the Rydberg blockade radius but also well beyond it, offering clear advantages over conventional blockade-based protocols.
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce the theoretical model and describe the physical mechanism of the proposed scheme. In Sec. III, we present a comprehensive performance assessment, including the dependence on the Rydberg interaction strength and the robustness of the protocol against Rydberg-state spontaneous emission, laser phase error, Rabi amplitude noise, and detuning errors. In Sec. IV, we analyze the application of our scheme to the realization of single-photon sources. In Sec. V, we discuss its implementation in a Rydberg superatom. In Sec. VI, we provide the corresponding energy level diagram and transitions. Finally, We conclude our work in Sec. VII.
@@ II.1 System Hamiltonian and Theoretical Framework
As illustrated in Fig. 1(a), we consider a cold ensemble of $^{87}\text{Rb}$ atoms confined in an optical dipole trap. The collective blockade radius is given by $R_{b}=[C_{6}/(\hbar\sqrt{\mathbb{N}}\Omega)]^{1/6}$. Here, $C_{6}$ is the van der Waals dispersion coefficient, and $\Omega$ denotes the effective Rabi frequency between the ground state and the Rydberg state. For simplicity, we adopt $\hbar=1$ in the subsequent analysis. The relevant atomic energy levels are shown in Fig. 1(b). In the interaction picture, the ground states $|g\rangle$ and $|e\rangle$ are coupled by a driving field with two-photon Rabi frequency $\omega$, described by the Hamiltonian
@@
Here, $U_{rr}$ denotes the Rydberg-mediated interaction that occurs when two atoms are simultaneously excited to the Rydberg state. This interaction is of the van der Waals type and can be expressed as $U_{rr}=C_{6}/d_{i,j}^{6}$, where $d_{i,j}$ is the distance between the $i$-th and $j$-th atoms. The operation is performed periodically over a total of $N$ cycles, with each cycle having a period of $T=t_{a}+t_{b}$. Each cycle consists of two successive stages: during the interval $0\leq t<t_{a}$, the dynamics are governed by the Hamiltonian $H_{a}$, while in the subsequent interval $t_{a}\leq t<t_{b}$, the system evolves under the Hamiltonian $H_{b}$. For convenience, the durations are chosen as $t_{a}=\pi/({\sqrt{\mathbb{N}}N\omega})$ and $t_{b}=4\pi/\Omega$, respectively, corresponding to a collective $\pi$ pulse on the $|g\rangle\leftrightarrow|e\rangle$ transition across all atoms and a $4\pi$ pulse on the $|e\rangle\leftrightarrow|r\rangle$ transition [76, 80]. The stroboscopic evolution of the system after each period is described by the Floquet evolution operator $U_{N}(T)$, which governs the dynamics of discrete time. The explicit expression for $U_{N}(T)$ is given as follows:
@@
$H_{z}$ is referred to as a Zeno Hamiltonian. $P_{n}$ represents the spectral projection of $U_{b}$ corresponding to the eigenvalue $\lambda_{n}$. For which an analytical expression is derived in Appendix A.
Figure 2: (a) Floquet quasienergy spectrum of the system. (b-c) Two representative regions extracted from panel (a), showing the comparison between the Floquet quasienergy spectrum and the effective coupling obtained from the unitary kicks Hamiltonian for different driving periods NN, in accordance with the quasienergy periodicity relation $\epsilon\equiv\epsilon+2n\pi/T$. (d) and (e) show population of the states for a two-atom system under $N=20$ and $N=30$ cycles, respectively. We set $\omega/2\pi=1$ MHz, $\Omega=5\omega$, and $U_{rr}=45\Omega$.
@@ II.2 Mechanism of Enhanced Ground-State Interaction and Ground-State Blockade
From the perspective of unitary kicks, because the system is initialized in $|gg\rangle$, the stroboscopic evolution is confined to the $H_{z}$ subspace, and the transition to the doubly excited state $|ee\rangle$ is strongly suppressed. The effective Hamiltonian can therefore be written as $H^{\text{eff}}_{Z}=H_{z}t_{a}/(t_{a}+t_{b})$, where $H_{z}=\frac{\sqrt{2}\omega}{2}(|gg\rangle\langle W_{2}|+\mathrm{H.c.})$. In other words, the unitary kicks effectively decouple $|ee\rangle$ from the dynamics. In fact, from a numerical perspective, the above dynamical process is independent of the number of driving periods $N$. For any value of $N$, the numerical results can be well explained by the effective Hamiltonian derived from the Floquet theorem. The corresponding effective Hamiltonian can be written as $H_{\mathrm{eff}}=i\mathrm{ln}(e^{-iH^{(1)}_{b}t_{b}}e^{-iH^{(1)}_{a}t_{a}})/T$, which allows us to further verify the precision of the analytical results through numerical simulations (see Appendix B). To verify the validity of this approach, we further consider finite values of $N$ and compare the resulting quasienergy spectrum with the Floquet theorem, as shown in Fig. 2(a). Each curve represents the quasienergy corresponding to an eigenstate, which is expressed as a linear superposition of the same set of basis vectors with different weights for different values of $N$. It can be seen that, for small $N$, the five branches of the quasienergy spectrum lie close to each other. With increasing $N$, the structure of the energy level gradually opens up, with $E_{1}$ and $E_{5}$ shifting towards higher and lower energies, respectively, while $E_{2}$, $E_{3}$, and $E_{4}$ converge to zero.In the large-$N$ limit, the eigenstate structure exhibits a clear pattern: the eigenstates associated with $E_{1}$ and $E_{5}$ are primarily composed of $|ee\rangle$ and $|D_{2}\rangle$, and can be approximately written as $|\Psi_{1,5}\rangle\approx c_{1}|D_{2}\rangle\mp c_{2}|ee\rangle$; similarly, the eigenstates corresponding to $E_{2}$ and $E_{4}$ are dominated by $|gg\rangle$ and $|W_{2}\rangle$, expressed as $|\Psi_{2,4}\rangle\approx c_{3}|gg\rangle\mp c_{4}|W_{2}\rangle$, while the eigenstate associated with $E_{3}$ is essentially given by $|T_{2}\rangle$ and $|\Psi_{3}\rangle\approx|T_{2}\rangle$, with all other components suppressed below $10^{-2}$, rendering their influence on the dynamics negligible. As $N$ increases, the effective dimensionality of the eigenstates decreases significantly, and the spectral structure becomes more distinguishable. Reconstructing $H_{\mathrm{eff}}=\sum_{i}E_{i}|\Psi_{i}\rangle\langle\Psi_{i}|$ and projecting it onto the reduced subspace spanned by $\{|gg\rangle,|W_{2}\rangle,|ee\rangle,|D_{2}\rangle\}$, we identify the dominant effective couplings that govern the dynamics. In this approximation, the quasienergy branch $E_{1}$ is equivalent to the coupling between $|ee\rangle$ and $|D_{2}\rangle$ with strength $\epsilon$, and $E_{2}$ is equivalent to the coupling between $|gg\rangle$ and $|W_{2}\rangle$ with strength $\mu$. These effective couplings capture the essential structure of the quasienergy spectrum and provide a clear physical interpretation of the underlying dynamics.
@@
Furthermore, Figs. 2(b) and 2(c) present the analytical results of the unitary kicks and the numerical results corresponding to the quasienergy spectrum shown in Fig. 2(a). The green line in Fig. 2(b) represents the effective coupling strength $\epsilon$ between $|ee\rangle$ and $|D_{2}\rangle$, corresponding to the quasienergy branch $E_{1}$ shown in Fig. 2(a). The orange and blue curves in Fig. 2(b) represent the quasienergy periodicity relation $\epsilon\equiv\epsilon+2n\pi/T$ for different $n$, while the orange line with circular markers represents the analytical results obtained from the unitary kicks. Owing to the intrinsic periodicity of the Floquet quasienergies, the analytical results coincide with the $n=2$ branch of the Floquet spectrum under the $4\pi$-pulse condition. As $N$ increases, the $|ee\rangle\to|D_{2}\rangle$ transition becomes significantly stronger, analogous to the enhancement of dipole-dipole interactions in Rydberg atomic ensembles. The solid red line in Fig. 2(c) represents the effective coupling strength $\mu$ between $|gg\rangle$ and $|W_{2}\rangle$, corresponding to the quasienergy branch $E_{2}$ shown in Fig. 2(a), while the orange line with circular markers represents the analytical results obtained from the unitary kicks. The green line denotes the numerical result of the Floquet theorem for the effective coupling strength $\lambda$ associated with the transition between $|ee\rangle$ and $|W_{2}\rangle$. We can observe that as $N$ increases, $\lambda$ and $\mu$ gradually decrease. This indicates that when the time scale $t_{a}$ is small, there exists a strong coupling interaction between the $|ee\rangle$ and $|D_{2}\rangle$ states, while the coupling from the $|ee\rangle$ to the $|W_{2}\rangle$ state is relatively weak. This result is similar to electromagnetically induced transparency, which prevents the transition from $|gg\rangle$ to $|ee\rangle$, thus evolving only from $|gg\rangle$ to $|W_{2}\rangle$. Usually, ground-state interactions are relatively weak at micrometer-scale interatomic distances. In other words, our scheme indirectly enhances the interaction strength between ground states. In this paper, we use the QUTIP package in Python for numerical analysis [82, 83].
@@
A sufficiently strong interaction between ground-state atoms gives rise to a ground-state blockade, in which the strong mutual interactions effectively suppress multiple ground-state excitations, thereby enabling the preparation of entangled states within the ground-state manifold. As illustrated in Fig. 2(d), we perform numerical simulations of a two-atom system governed by Eq. (3) and Eq. (II.1) to examine the feasibility of realizing ground-state blockade. The blue and yellow lines represent the populations of the states $|W_{2}\rangle$ and $|gg\rangle$, respectively. The solid line and the line with circular markers represent the numerical evolution of Eq. (3) and the effective evolution of Eq. (II.1), respectively. The results clearly show that the system achieves a complete population transfer from $|gg\rangle$ to $|W_{2}\rangle$ after $20$ driving cycles, consistent for both numerical and effective solutions. These results confirm the validity of our theoretical analysis and demonstrate that ground-state blockade can be successfully realized through enhanced ground-state interactions using our scheme, achieving a final population of approximately $99.2\%$ in the $|W_{2}\rangle$ state. In addition, we simulate the dynamic evolution process of state $|gg\rangle$ and state $|W_{2}\rangle$ with the time cycles $N=30$, as shown in Fig. 2(e). The results show that the population of the target state reaches about $99.6\%$, confirming that the effective evolution remains valid even for large $N$. To ensure simulation precision, all subsequent numerical simulations employ the numerical solution of Eq. (3) to validate the realization of the ground-state blockade. In the main text, we only show the stroboscopic times for the dynamic process, while Appendix C provides a comprehensive demonstration of the continuous-time evolution.

### v2

@@ I Introduction
Ground-state neutral atoms exhibit long coherence times and excellent isolation from environmental noise, making them promising candidates for scalable quantum technologies [1, 2, 3, 4, 5, 6, 7]. Meanwhile, coupling ground states to Rydberg levels offers a complementary advantage: strong, long-range interactions via van der Waals (vdW) or dipole-dipole interactions, which enable fast entangling operations through the Rydberg blockade effect [8, 9, 10, 11, 12, 13]. The Rydberg blockade provides a powerful mechanism for engineering strong, controllable interactions and creating collective quantum states in which a single excitation is coherently shared throughout the atomic ensemble [14, 15, 16]. Such collective control has opened broad application prospects in quantum optics [17, 18, 19, 20, 21, 22, 23], quantum computing [24, 25, 26], quantum communication [27, 28, 29, 30, 31], and single-photon sources [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. With the development of experimental technology, the coherent manipulation of Rydberg atomic ensembles is driving a deeper understanding of collective quantum dynamics and enabling new opportunities for scalable and high-performance quantum technologies [43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
The finite lifetime and strong decoherence effects inherent to Rydberg states present significant challenges for high-fidelity quantum operations. This naturally raises the question: Can sufficiently strong and controllable interactions be engineered while atoms remain in their ground state? In this context, finding ways to achieve effective interactions without exciting high-energy states is especially important. Just as the Rydberg blockade allows only a single collective excitation in Rydberg states, strong interactions between ground states can also give rise to a ground-state blockade mechanism [53, 54, 55, 56, 57, 58]. Consequently, the system is deterministically prepared in the target symmetric collective state in the ground-state manifold. This symmetric collective state is of great significance for photon storage [59, 60], single-photon sources [61, 62, 63, 64, 65], quantum state transfer [66, 67], and other quantum information technologies [68, 69]. Compared to the complexity and strong interactions inherent in Rydberg states, the symmetric collective state of ground states provides a more precise and stable quantum platform, making it particularly promising for applications in quantum computing and quantum simulation [70, 71, 72, 73].
@@
In this work, we utilize Floquet stroboscopic dynamics to prepare a $W$ state within the ground-state manifold of a Rydberg atomic ensemble [74, 75, 76, 77, 78, 79, 80, 81, 82]. The system evolves from the ground state $|G\rangle=|g_{1}g_{2}...g_{\mathbb{N}}\rangle$ to the symmetric collective state $|W_{\mathbb{N}}\rangle={1/\sqrt{\mathbb{N}}\textstyle\sum_{i=1}^{\mathbb{N}}}|g_{1}g_{2}...e_{i}...g_{\mathbb{N}}\rangle$, while the double excitations remain unpopulated at stroboscopic time points. This indirectly induces an effective interaction between ground-state neutral atoms, leading to a strong blockade effect that enables high-fidelity preparation of the $W$ state. Finally, we analyze the application of the scheme to single-photon generation. Our approach remains effective not only within the Rydberg blockade radius but also well beyond it, offering clear advantages over conventional blockade-based protocols.
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce the theoretical model and describe the physical mechanism of the proposed scheme. In Sec. III, we present a comprehensive performance assessment, including the dependence on the Rydberg interaction strength and the robustness of the protocol against various experimentally relevant imperfections, such as Rydberg-state decay (including blackbody radiation), laser phase and amplitude noise, detuning errors, and Doppler-induced collective dephasing. We also compare our scheme with several representative approaches. In Sec. IV, we analyze the application of our scheme to the realization of single-photon sources. In Sec. V, we discuss its implementation in a Rydberg superatom. In Sec. VI, we provide the corresponding energy level diagram and transitions. Finally, We conclude our work in Sec. VII.
@@ II.1 System Hamiltonian and Theoretical Framework
As illustrated in Fig. 1(a), we consider a cold ensemble of $^{87}\text{Rb}$ atoms confined in an optical dipole trap. The collective blockade radius is given by $R_{b}=[|C_{6}|/(\hbar\sqrt{\mathbb{N}}\Omega)]^{1/6}$. Here, $C_{6}$ is the vdW dispersion coefficient, and $\Omega$ denotes the effective Rabi frequency between the ground state and the Rydberg state. For simplicity, we adopt $\hbar=1$ in the subsequent analysis. The relevant atomic energy levels are shown in Fig. 1(b). In the interaction picture,the ground states $|g\rangle$ and $|e\rangle$ are coupled by a coherent driving field, which can be implemented either as a direct microwave transition or as an effective two-photon Raman coupling, with Rabi frequency $\omega$, described by the Hamiltonian
@@
Here, $U_{rr}$ denotes the Rydberg-mediated interaction that occurs when two atoms are simultaneously excited to the Rydberg state. This interaction is of the vdW type and can be expressed as $U_{rr}=-C_{6}/d_{i,j}^{6}$, where $d_{i,j}$ is the distance between the $i$-th and $j$-th atoms. The operation is performed periodically over a total of $N$ cycles, with each cycle having a period of $T=t_{a}+t_{b}$. Each cycle consists of two successive stages: during the interval $0\leq t<t_{a}$, the dynamics are governed by the Hamiltonian $H_{a}$, while in the subsequent interval $t_{a}\leq t<t_{b}$, the system evolves under the Hamiltonian $H_{b}$. For convenience, the durations are chosen as $t_{a}=\pi/({\sqrt{\mathbb{N}}N\omega})$ and $t_{b}=4\pi/\Omega$, respectively, corresponding to a collective $\pi$ pulse on the $|g\rangle\leftrightarrow|e\rangle$ transition across all atoms and a $4\pi$ pulse on the $|e\rangle\leftrightarrow|r\rangle$ transition [78, 82]. The stroboscopic evolution of the system after each period is described by the Floquet evolution operator $U_{N}(T)$, which governs the dynamics of discrete time. The explicit expression for $U_{N}(T)$ is given as follows:
@@
$H_{z}$ is referred to as the Zeno Hamiltonian. $P_{n}$ represents the spectral projection of $U_{b}$ corresponding to the eigenvalue $\lambda_{n}$. For which an analytical expression is derived in Appendix A.
Figure 2: (a) Floquet quasienergy spectrum of the system. (b-c) Two representative regions extracted from panel (a), showing the comparison between the Floquet quasienergy spectrum and the effective coupling obtained from the unitary kicks Hamiltonian for different driving periods NN, in accordance with the quasienergy periodicity relation $\epsilon\equiv\epsilon+2n\pi/T$. (d) and (e) show population of the states for a two-atom system under $N=30$ and $N=20$ cycles, respectively. We set $\omega/2\pi=1$ MHz, $\Omega=5\omega$, and $U_{rr}=45\Omega$.
@@ II.2 Mechanism of Enhanced Ground-State Interaction and Ground-State Blockade
From the perspective of unitary kicks, because the system is initialized in $|gg\rangle$, the stroboscopic evolution is confined to the $H_{z}$ subspace, and the transition to the doubly excited state $|ee\rangle$ is strongly suppressed. The effective Hamiltonian can therefore be written as $H^{\text{eff}}_{Z}=H_{z}t_{a}/(t_{a}+t_{b})$, where $H_{z}=\frac{\sqrt{2}\omega}{2}(|gg\rangle\langle W_{2}|+\mathrm{H.c.})$. In other words, the unitary kicks effectively decouple $|ee\rangle$ from the dynamics. In fact, from a numerical perspective, the above dynamical process is independent of the number of driving periods $N$. For any value of $N$, the numerical results can be well explained by the effective Hamiltonian derived from the Floquet theorem. The corresponding effective Hamiltonian can be written as $H_{\mathrm{eff}}=i\mathrm{ln}(e^{-iH^{(1)}_{b}t_{b}}e^{-iH^{(1)}_{a}t_{a}})/T$, which allows us to further verify the precision of the analytical results through numerical simulations (see Appendix B). To verify the validity of this approach, we further consider finite values of $N$ and compare the resulting quasienergy spectrum with the Floquet theorem, as shown in Fig. 2(a). Each curve represents the quasienergy corresponding to an eigenstate, which is expressed as a linear superposition of the same set of basis vectors with different weights for different values of $N$. It can be seen that, for small $N$, the five branches of the quasienergy spectrum are close to each other. With increasing $N$, the structure of the energy gap gradually opens up, with $E_{1}$ and $E_{5}$ shifting towards higher and lower energies, respectively, while $E_{2}$, $E_{3}$, and $E_{4}$ converge to zero. In the large-$N$ limit, the eigenstate structure exhibits a clear pattern: the eigenstates associated with $E_{1}$ and $E_{5}$ are primarily composed of $|ee\rangle$ and $|D_{2}\rangle$, and can be approximately written as $|\Psi_{1,5}\rangle\approx c_{1}|D_{2}\rangle\mp c_{2}|ee\rangle$; similarly, the eigenstates corresponding to $E_{2}$ and $E_{4}$ are dominated by $|gg\rangle$ and $|W_{2}\rangle$, expressed as $|\Psi_{2,4}\rangle\approx c_{3}|gg\rangle\mp c_{4}|W_{2}\rangle$, while the eigenstate associated with $E_{3}$ is essentially given by $|T_{2}\rangle$ and $|\Psi_{3}\rangle\approx|T_{2}\rangle$, with all other components suppressed below $10^{-2}$, rendering their influence on the dynamics negligible. As $N$ increases, the effective dimensionality of the eigenstates decreases significantly, and the spectral structure becomes more distinguishable. Reconstructing $H_{\mathrm{eff}}=\sum_{i}E_{i}|\Psi_{i}\rangle\langle\Psi_{i}|$ and projecting it onto the reduced subspace spanned by $\{|gg\rangle,|W_{2}\rangle,|ee\rangle,|D_{2}\rangle\}$, we identify the dominant effective couplings that govern the dynamics. In this approximation, the quasienergy branch $E_{1}$ is equivalent to the coupling between $|ee\rangle$ and $|D_{2}\rangle$ with strength $\epsilon$, and $E_{2}$ is equivalent to the coupling between $|gg\rangle$ and $|W_{2}\rangle$ with strength $\mu$. These effective couplings capture the essential structure of the quasienergy spectrum and provide a clear physical interpretation of the underlying dynamics.
@@
Furthermore, Figs. 2(b) and 2(c) present the analytical results of the unitary kicks and the numerical results corresponding to the quasienergy spectrum shown in Fig. 2(a). The green line in Fig. 2(b) represents the effective coupling strength $\epsilon$ between $|ee\rangle$ and $|D_{2}\rangle$, corresponding to the quasienergy branch $E_{1}$ shown in Fig. 2(a). The orange and blue curves in Fig. 2(b) represent the quasienergy periodicity relation $\epsilon\equiv\epsilon+2n\pi/T$ for different $n$, where $n$ is an integer indexing the periodic Floquet branches [84, 85], while the orange line with circular markers represents the analytical results obtained from the unitary kicks. Owing to the intrinsic periodicity of the Floquet quasienergies, the analytical results coincide with the $n=2$ branch of the Floquet spectrum under the $4\pi$-pulse condition. As $N$ increases, the $|ee\rangle\to|D_{2}\rangle$ transition becomes significantly stronger, analogous to the enhancement of dipole-dipole interactions in Rydberg atomic ensembles. The solid red line in Fig. 2(c) represents the effective coupling strength $\mu$ between $|gg\rangle$ and $|W_{2}\rangle$, corresponding to the quasienergy branch $E_{2}$ shown in Fig. 2(a), while the orange line with circular markers represents the analytical results obtained from the unitary kicks. The green line denotes the numerical result of the Floquet theorem for the effective coupling strength $\lambda$ associated with the transition between $|ee\rangle$ and $|W_{2}\rangle$. We can observe that as $N$ increases, $\lambda$ and $\mu$ gradually decrease. This indicates that when the time scale $t_{a}$ is small, there exists a strong coupling interaction between the $|ee\rangle$ and $|D_{2}\rangle$ states, while the coupling from the $|ee\rangle$ to the $|W_{2}\rangle$ state is relatively weak. This result is similar to Autler-Townes splitting, which prevents the transition from $|gg\rangle$ to $|ee\rangle$, thus evolving only from $|gg\rangle$ to $|W_{2}\rangle$. Usually, ground-state interactions are relatively weak at micrometer-scale interatomic distances. In other words, our scheme indirectly enhances the interaction strength between ground states. This induced strong coupling gives rise to a ground-state blockade, in which the simultaneous population of multiple atoms in the $|e\rangle$ state is effectively suppressed, thereby enabling the preparation of entangled states within the ground-state manifold.
@@
Having analyzed the static quasienergy spectrum, we now demonstrate the blockade effect through dynamical simulations. As shown in Fig. 2(d), we numerically simulate a two-atom system governed by Eqs. (3) and (II.1) to assess the feasibility of realizing ground-state blockade. The blue and yellow curves denote the populations of the states $|W_{2}\rangle$ and $|gg\rangle$, respectively. Solid lines correspond to the full numerical evolution under Eq. (3), while curves with circular markers represent the effective dynamics described by Eq. (II.1). We first consider the dynamical evolution over $N=30$ driving cycles. In this case, the population of the target state $|W_{2}\rangle$ reaches approximately $99.6\%$, demonstrating excellent agreement between the full and effective descriptions and confirming the validity of the effective evolution for sufficiently large $N$. Figure 2(e) presents the results for a reduced number of cycles, $N=20$, where the scheme remains robust and achieves nearly complete population transfer from $|gg\rangle$ to $|W_{2}\rangle$. Both the numerical and effective solutions exhibit consistent behavior, yielding a final $|W_{2}\rangle$ population of about $99.2\%$. These results validate our theoretical analysis and confirm that ground-state blockade can be efficiently realized via enhanced ground-state interactions within the proposed scheme. To ensure numerical accuracy, all subsequent simulations are based on the full numerical solution of Eq. (3). In the main text, we present only the stroboscopic dynamics, while Appendix C provides the corresponding continuous-time evolution. All numerical calculations are performed using the QuTiP package in Python [86, 87].